# Frontier-Based Exploration for Multi-Robot Rendezvous in Communication-Restricted Unknown Environments

### v2

@@ II State of the art
Methods for multi-robot rendezvous typically leverage pre-determined coordination strategies among robots (pre-established rendezvous areas, biases applied to strategies to search for others, etc.) or execute online search strategies in a commonly known map to meet others. One distinction among existing approaches can be made between symmetric and asymmetric rendezvous. In symmetric rendezvous, all robots have the same role in seeking a meeting with others. In the asymmetric case, instead, some robots can be explorers, while others are relays, i.e., they have to meet and transfer knowledge acquired by other robots to each other or to a base station. Another distinction is between intentional rendezvous and accidental rendezvous. The former ones happen when the meeting is already scheduled among agents following some kind of coordination. The latter ones take place without a pre-arrangement. Other approaches undertake an offline study of the problem, by deriving theoretical properties of the optimal solution from the specific geometrical features of the environment [11]. Several of these approaches are not directly applicable to our scenario, where the target environment and starting locations of the MRS are unknown, and communication is restricted.
@@
The work of [17] presents a framework to perform exploration, search, and rendezvous in a real-world deployment, by relying on the content of virtual stigmergy [18]. Similarly to us, the robots are initially distributed in the environment and communicate only when are in a line of sight. However, the goal of this work is significantly different from ours, as it aims to perform periodic rendezvous at a pre-arranged time to share the information acquired during single-robot search tasks. After that, the MRS sets up a relay tree to connect the team to a base, sharing the position of the targets identified.
@@ III-B Adapting Frontier-Based Exploration for Rendezvous
    rank frontiers according to one or more criteria aimed at determining the effectiveness of continuing the exploration by traveling there, and selecting the best one;
@@
A key part of our method is to augment the set of frontiers $F$ with additional "virtual" frontiers that do not adhere to the standard definition but instead represent regions of interest for rendezvous. These frontiers, which we denote with the set $\bar{F}$, are obtained by using an information decay mechanism on the exploration trace. More precisely, at time $t$, we consider a portion of the exploration trace that is limited to the sub-trajectory linking the last $k$ poses, $P^{k}_{i}=\{p^{i}_{t-k+1},\ldots,p^{i}_{t}\}$, which start at the robot’s current position and are evenly spaced from there along the past visited locations. Each pose is characterized by a timestamp indicating when it was obtained; poses are removed after a time $T$ has passed after their acquisition, hence partitioning the trace into an active head and a vanished tail. Each time a pose and the corresponding footprint on the trace are removed, a new virtual frontier $\bar{f}\in\bar{F}$ is created as the contour of the difference between the areas before and after the removal. A visual intuition of this process and its result is depicted in Fig. 3.
@@ III-C Online cluster formation
Figure 4: (left) Cluster formation; (right) Example of two robots that, after they meet, share their exploration trace. The red robot’s frontier 2 is deleted after the meeting as it overlaps the exploration trace of the blue robot.
@@ IV-A Experimental Setting
We indicate as success rate $R$ as the percentage of runs concluded with a rendezvous, and the time $t_{i}$ as the time required for $i\in\{1,\ldots,N$} robots to do a (partial) rendezvous. Our average results do not include runs where one or more robots experienced a permanent failure during navigation (e.g., due to a collision or because incapable of maneuvering away from a too-close obstacle). In our realistic setup, these kinds of events are fairly frequent, but in the majority of cases, robots managed to resolve the impasse by triggering a recovery behavior on the navigation stack.
In the following, we provide a significant selection of the obtained results. The used code and the full set of experimental results are available in a public repository111 https://aislabunimi.github.io/fberendezvous/.
@@ IV-B Results
The results are reported in Table I. We first discuss the results obtained in Ring and Office, using 3, 5, and 8 robots. The first interesting remark that emerges is that a frontier-based exploration strategy can be used, with fair results, to perform a rendezvous in an initially unknown environment. At the same time, the use of our method significantly improves the chances of episodic rendezvous, as can be seen by the gain in terms of $\Delta_{t}$. This can be seen from the different paths resulting from the two methods, as shown in Fig. 1 and Fig. 5. The improvement is more evident when the number of robots is small, and the environment is large, i.e., when there are a few robots distributed across a large map. While our method allowed the whole team of robots to perform a rendezvous in all of the $60$ runs performed, with $R=1$, there is a non-negligible chance that robots are not able to perform a rendezvous if a standard frontier-based exploration method FBE is used. This usually happens when, by chance, robots in the early stages of exploration choose different directions; during the remainder of the exploration run, the robots switched their location without meeting the other members of the MRS. An example of this can be seen in Fig. 1(right). The use of our method, which encourages the robot to backtrack on previously explored positions, is effective in facilitating rendezvous. At the same time, the performance is far more stable than those of FBE, with a lower standard deviation. These findings are reported in Fig. 6, which shows the result in terms of time $t$ for 3 to 8 robots in Ring. When more robots are used, the chances of having episodic rendezvous are higher, and thus the differences between the two methods are less evident. In that case, the standard FBE can be effectively used also to perform rendezvous.
FBR FBE Environment # $t$ (s) $\sigma_{t}$ $R$ $t$ (s) $\sigma_{t}$ $R$ $\Delta_{t}$ Ring 3 1623.13 978.18 1 2691.64 1557.18 0.6 0.66 5 2000 936.3 1 2846.06 1246.36 0.7 0.42 8 1844.08 488.16 1 2003.4 782.05 1 0.09 Office 3 727.46 372.16 1 851.11 791.46 0.9 0.17 5 1358.55 718.19 1 1478.21 594.59 8.8 0.02 8 1309.45 437.5 1 1251.54 762.47 1 -0.04 Intel 3 580.13 256.63 1 688.19 553.28 0.9 0.19 Campus-1 3 1070.29 487.82 1 1591.7 1130.81 1 0.49 Campus-2 3 818.09 503.57 1 1370.68 959.97 1 0.68
TABLE I: Results of FBR compared with FBE in all of the environments $t$ is in seconds, $R$ is in $[0,1]$. With $\Delta_{t}$ we indicate the $\Delta_{t}=t_{\texttt{FBR}}/t_{\texttt{FBE}}-1$.
@@
Figure 6: (left) The average time $t$ (and standard deviation) to complete the rendezvous changing the maximum size of the cluster in the environment ring when an MRS of $5$ robots is used; note how, in this environment the $R=0.7$ for FBE. (right) The average time $t$ (and standard deviation) required for a run in ring when a different number of robots is used.

### v3

@@ II State of the art
Methods for multi-robot rendezvous typically leverage pre-determined coordination strategies among robots (pre-established rendezvous areas, biases applied to strategies to search for others, etc.) or execute online search strategies in a commonly known map to meet others. One distinction among existing approaches can be made between symmetric and asymmetric rendezvous. In symmetric rendezvous, all robots have the same role in seeking a meeting with others. In the asymmetric case, instead, some robots can be explorers, while others are relays, i.e., they have to meet and transfer knowledge acquired by other robots to each other or to a base station. Another distinction is between intentional rendezvous and accidental rendezvous. The former ones happen when the meeting is already scheduled among agents following some kind of coordination. In the latter ones the robots know that they need to meet but no time and place is pre-arranged. Other approaches undertake an offline study of the problem, by deriving theoretical properties of the optimal solution from the specific geometrical features of the environment [11]. Several of these approaches are not directly applicable to our scenario, where the target environment and starting locations of the MRS are unknown, and communication is restricted.
@@
The work of [17] presents a framework similar to ours where to perform exploration, search, and rendezvous in a real-world deployment with line-of-sight communication, by relying on the content of virtual stigmergy [18]. However, the goal of such a work is significantly different from ours, as it aims to perform periodic rendezvous at a pre-arranged time to share the information acquired during single-robot search tasks. After that, the MRS sets up a relay tree to connect the team to a base, sharing the position of the targets identified.
@@ III-B Adapting Frontier-Based Exploration for Rendezvous
    rank frontiers according to one or more criteria (e.g., the distance from the robot) aimed at determining the effectiveness of continuing the exploration by traveling there, and selecting the best one;
@@
A key part of our method is to augment the set of frontiers $F$ with additional "virtual" frontiers that do not adhere to the standard definition but instead represent regions of interest for rendezvous. These frontiers, which we denote with the set $\bar{F}$, are obtained by using an information decay mechanism on the exploration trace. More precisely, at time $t$, we consider a portion of the exploration trace that is limited to the sub-trajectory linking the last $k$ poses. Denoting with $p^{i}_{t}$ the pose of robot $i$ at time step $t$, the trace is represented by the set $P^{i}_{k}=\{p^{i}_{t-k+1},\ldots,p^{i}_{t}\}$. The trace portion starts at the robot’s current position and are evenly spaced from there along the past visited locations. Each pose is characterized by a timestamp indicating when it was obtained; poses are removed after a time $T$ has passed after their acquisition, hence partitioning the trace into an active head and a vanishing tail. Each time a pose and the corresponding footprint on the trace are removed, a new virtual frontier $\bar{f}\in\bar{F}$ is created as the contour of the difference between the areas before and after the removal. A visual intuition of this process and its result is depicted in Fig. 3.
@@ III-C Online cluster formation
Figure 4: (left) Cluster formation; (right) Two robots meet and share their exploration trace. The frontiers are numbered. The red robot’s frontier (2) is deleted as it overlaps with the exploration trace of the blue robot.
@@ IV-A Experimental Setting
We indicate as success rate $R$ as the percentage of runs concluded with a rendezvous, and the time $t_{i}$ as the time required for $i\in\{1,\ldots,N$} robots to do a rendezvous. Our average results do not include runs where one or more robots experienced a permanent failure during navigation (e.g., due to a collision or because incapable of maneuvering away from a too-close obstacle). In our realistic setup, these kinds of events are fairly frequent, but in the majority of cases, robots managed to resolve the impasse by triggering a recovery behavior on the navigation stack.
In the following, we provide a significant selection of the obtained results. The used code and the full set of experimental results are available in a public repository [^1] https://aislab.di.unimi.it/research/fberendezvous/.
@@ IV-B Results
The results are reported in Table I. We first discuss the results obtained in Ring and Office, using 3, 5, and 8 robots. The first interesting remark that emerges is that a frontier-based exploration strategy can be used, with fair results, to perform a rendezvous in an initially unknown environment. At the same time, the use of our method significantly improves the chances of episodic rendezvous, as can be seen by the gain in terms of $\Delta_{t}=(t_{\texttt{FBR}}-t_{\texttt{FBE}})/t_{\texttt{FBE}}$, where $t_{\texttt{FBR}}$ and $t_{\texttt{FBE}}$ are the times required by FBR and FBE, respectively. This can be seen from the different paths resulting from the two methods, as shown in Fig. 1 and Fig. 5. The improvement is more evident when the number of robots is small, and the environment is large, i.e., when there are a few robots distributed across a large map. While our method allowed the whole team of robots to perform a rendezvous in all of the $60$ runs of this type, with $R=1$, there is a non-negligible chance that robots are not able to perform a rendezvous if a standard frontier-based exploration method FBE is used. This usually happens when, by chance, robots in the early stages of exploration choose different directions; during the remainder of the exploration run, the robots switched their location without meeting the other members of the MRS. An example of this can be seen in Fig. 1(right). The use of our method, which encourages the robot to backtrack on previously explored positions, is effective in facilitating rendezvous. At the same time, the performance is far more stable than those of FBE, with a lower standard deviation. These findings are reported in Fig. 6, which shows the result in terms of time $t$ for 3 to 8 robots in Ring. When more robots are used, the chances of having episodic rendezvous are higher, and thus the differences between the two methods are less evident. In that case, the standard FBE can be effectively used also to perform rendezvous.
FBR FBE Environment $m$ $t$ (s) $\sigma_{t}$ $R$ $t$ (s) $\sigma_{t}$ $R$ $\Delta_{t}$ Ring 3 1623.13 978.18 1 2691.64 1557.18 0.6 0.66 5 2000 936.3 1 2846.06 1246.36 0.7 0.42 8 1844.08 488.16 1 2003.4 782.05 1 0.09 Office 3 727.46 372.16 1 851.11 791.46 0.9 0.17 5 1358.55 718.19 1 1478.21 594.59 8.8 0.02 8 1309.45 437.5 1 1251.54 762.47 1 -0.04 Intel 3 580.13 256.63 1 688.19 553.28 0.9 0.19 Campus-1 3 1070.29 487.82 1 1591.7 1130.81 1 0.49 Campus-2 3 818.09 503.57 1 1370.68 959.97 1 0.68
TABLE I: Results of FBR compared with FBE in all of the environments $t$ is in seconds, $R$ is in $[0,1]$.
@@
Figure 6: (left) The average time $t$ (and standard deviation) to complete the rendezvous as the maximum size of the cluster changes in the environment ring when an MRS of $5$ robots is used; note how, in this environment the $R=0.7$ for FBE. (right) The average time $t$ (and standard deviation) required for a run in ring when a different number of robots is used.